# Human2Robot: Learning Robot Actions from Paired Human-Robot Videos

### v2

@@ Abstract
Distilling knowledge from human demonstrations is a promising way for robots to learn and act. Existing work often overlooks the differences between humans and robots, producing unsatisfactory results. In this paper, we study how perfectly aligned human-robot pairs benefit robot learning. Capitalizing on VR-based teleportation, we introduce H&R, a third-person dataset with 2,600 episodes, each of which captures the fine-grained correspondence between human hand and robot gripper. Inspired by the recent success of diffusion models, we introduce Human2Robot, an end-to-end diffusion framework that formulates learning from human demonstration as a generative task. Human2Robot fully explores temporal dynamics in human videos to generate robot videos and predict actions at the same time. Through comprehensive evaluations of 4 carefully selected tasks in real-world settings, we demonstrate that Human2Robot can not only generate high-quality robot videos but also excels in seen tasks and generalizing to different positions, unseen appearances, novel instances, and even new backgrounds and task types.
@@ 1 Introduction
Figure 1: Human2Robot: An end-to-end generative framework that simultaneously predicts robotic videos and actions.
The ability to observe others, whether people or animals, and learn skills to solve new tasks is a key reason humans can handle a wide range of complex challenges. To develop robots capable of assisting with various everyday problems, they must also acquire this skill, particularly the ability to learn directly from humans. This motivates a plethora of work studying how to effectively learn from human demonstrations [5, 56, 38, 60, 4, 53]. However, due to the discrepancies between humans and robots, existing work [60, 38, 71] struggles to transfer learned representations from humans to robots, or simply enforce consistent outputs between humans and robots [4] without considering motion trajectories.
In this paper, we advocate that fine-grained alignment between humans and robots is of great significance. Coarse-grained data provides limited clues for key points and cannot capture detailed information, such as different movement trajectories or grasping positions. While annotating detailed correspondences between humans and robots is feasible [29, 34], it is time-consuming and labor-intensive. Fortunately, virtual reality(VR)-based teleoperation [28, 15, 43, 13] offers a promising way to collect synchronized actions of humans and robots, making it efficient and affordable. This motivates us to build upon VR-based teleoperation for studying paired data that can improve robot learning. In particular, we focus on obtaining perfectly aligned data between humans and robots as a testbed for learning from human demonstrations.
To this end, we introduce a novel data collection pipeline for gathering paired videos of human hands and robotic arms, aligned frame by frame. Unlike previous work that solely collects demonstration data, our method requires minimal additional setup, equipment, or time, making it exceptionally easy to implement and facilitating efficient and consistent data collection. We alse introduce the H&R dataset, a third-person dataset featuring perfectly aligned videos of human hands and robotic arms, which includes 8 fundamental tasks and 6 long-horizon tasks, totalling 2,600 episodes. Our goal is to train a model that mimics human actions and performs corresponding tasks based on human demonstrations. This requires the robot to understand how its end effector is aligned with human hands, which is readily available in our dataset, but more importantly, has the ability to act correspondingly. Therefore, the robot needs to predict what actions to take in the future conditioned on human motions as well as historical clues. This is similar in spirit to video generation tasks, that predict future frames (actions) based on various cues [14, 23, 24, 25, 37, 41]. As a result, we build upon video generation models to mimic human actions.
In light of this, we introduce Human2Robot, an end-to-end generative framework that predicts robotic videos and actions to take simultaneously. In particular, Human2Robot is built upon a pretrained Stable Diffusion model, including a spatial UNet for feature extraction, two behavior extractors for motion and position encoding, a spatial-temporal UNet that leverages temporal clues for frame and action prediction. As a result, Human2Robot is not only able to generate high-quality robotic videos, which facilitating the expansion of robot-conditioned policies to human-conditioned ones, but can also predict corresponding action trajectory required to perform the task. With carefully designed training strategies, Human2Robot not only excels on seen tasks but also generalizes to different positions, unseen appearances, instances and even new backgrounds and task types without additional training. Furthermore, we introduce a KNN-based approach, enabling Human2Robot to perform seen tasks even without human demonstrations, providing a flexible and scalable solution to robotic learning from demonstrations.
@@
  • We introduce a novel low-cost, low-requirement, high-efficiency, and highly scalable pipeline for collecting perfectly aligned data of humans and robots at the frame level. Additionally, we present H&R, the first dataset featuring perfectly aligned videos of human hands and robotic arms across a variety of tasks, enabling high-fidelity learning from human demonstrations.
  • We introduce Human2Robot, an end-to-end generative framework built upon a diffusion-based model, capable of generating robotic videos from human videos while simultaneously predicting the corresponding robotic action trajectories. This approach combines state-of-the-art video generation techniques with multi-task action prediction, facilitating accurate task execution.
  • We propose a KNN+Human2Robot method, which integrates KNN for task prediction, enabling to perform tasks even without human hand videos as input. This further enhances the scalability and flexibility of the system.
  • Human2Robot can generate high-quality robotic arm videos, which facilitates the expansion of robot-conditioned policy to human-conditioned policy, while also excelling in four carefully selected tasks, even with variations in positions, appearances, instances, backgrounds and different task types.
@@ 2 Related Work
Teleoperation. Researchers have developed various forms of teleoperation approaches for data collection [33, 21, 52, 43, 55, 2, 3, 45, 19, 72, 17, 1]. However, these approaches may entail specific device requirements. Recently, VR-based methodologies [28, 15] have attracted considerable attention due to their cost-effectiveness, efficiency and versatility. However, these methods often overlook the importance of the paired visual information. We leverage this perception to capture perfectly aligned videos of humans and robots, which are crucial for robotic imitation.
Learning from Human Videos. Researchers have recently been attempting to leverage existing human-centric video datasets to enhance robot policy learning [34, 54, 10, 51, 71]. Researchers propose to learn representations from human videos to assist in task execution [65, 60, 36]. However, these approaches need substantial prior knowledge and struggle to transfer to robots [38, 5, 4]. Meanwhile, some human-video-conditioned methods  [7, 29, 69] have focused on aligning representations across human and robot videos, they are neither efficient nor capable of generalization. In contrast, our approach utilizes paired data and diffusion models to achieve strong generalization capabilities.
Diffusion Models for Video Generation. Due to the remarkable performance of diffusion models in image generation [46, 47, 40, 49, 6], many approaches have explored their potential for video generation [26, 16, 8, 64, 32, 67, 57, 58, 59]. Moreover, conditionally controlled video generation [39, 30, 61, 27, 66] has broader and more practical application scenarios. In recent work, several approaches [9, 63, 7] have utilized generative diffusion models to aid policy learning for robots. However, these methods primarily generate videos from a conventional human perspective, often overlooking the distributional differences between human and robotic actions. In contrast, we introduce a novel approach by applying video diffusion models to robotic scenarios, utilizing a generative model conditioned on human behaviors to guide robotic action outputs.
Diffusion Models for Visual Understanding. The diffusion model not only performs impressively in generative domains but also shows promising applications in visual understanding tasks [70, 50, 35, 18]. Some notable approaches involve using diffusion models for object detection [11], image segmentation [68], and visual representation learning [70, 12]. The approach most similar to ours, GenRec [62] proposes a unified video diffusion model that jointly optimizes video generation and video classification, exhibiting the effectiveness of the diffusion model in video understanding. Our approach is designed for robotic tasks, constructing a human-to-robot video generative model that can simultaneously predict actions for manipulation tasks.
@@ 3 H&R Dataset
3.1 Data Collection Pipeline
Our system setup is straightforward, consisting of a Meta Quest 3 VR headset, a 6-DoF xArm robotic arm with a xArm gripper, and two Intel Realsense D435 cameras capturing images of the human hand and robotic arm from the same viewpoint.
The dual-camera setup ensures capturing the synchronized movements of both the human and robotic arms. With the integration of the hardware components, our system aims to achieve a critical objective: Producing a paired dataset where each data point reflects a seamless synchronization of human hand action and robotic arm response, captured synchronously from identical viewpoints.
Towards this goal, we build our H&R system focusing on developing software and establishing data collection protocols. We build the control system with OpenTeach [28]. When the user puts on the headset and performs several hand operations, the VR headset sends the coordinates of the hand movements of the user to the robotic arm and achieves synchronized actions.
Achieving this functionality requires addressing issues such as coordinate alignment, precision, and jitter. The details of the implementation are provided in the Supplementary materials B.2.
@@ 3.2 Statistics of H&R Dataset
With our designed system and data collection pipeline, we propose our H&R dataset, the first dataset featuring paired human and robotic videos. Human events vary from 8 fundamental tasks and 6 long-horizon tasks. The whole dataset includes 2,600 episodes, and each episode contains frames ranging from 300 to 600. Details can be found in Supplementary materials C.
To the best of our knowledge, the H&R dataset is the first video dataset that ensures perfectly aligned video between human and robot.
Figure 2: Data demonstration. The image above shows a sample from our dataset, recorded simultaneously by two cameras—one for the human and one for the robot—while the user wears a VR headset to perform the operation using our pipeline.
@@ 4 Human2Robot
Figure 3: Architecture overview of Human2Robot. Our approach consists of two training stages. In the first stage, only Behavior Extractors and Spatial Layer in both Spatial UNet and Spatial-Temporal UNet are trainable. In the second stage, we only train Temporal Layer and Policy Head. The Pretrained Encoder and Decoder are frozen all the time. In inference, we have two different modes: the KNN-Based mode can complete seen tasks without hand video, while the Manual Guided mode can complete both seen and unseen tasks. Proprio refers to the proprioception of the robot.
Our goal is to develop a model capable of learning from human demonstrations to predict future robot observations and corresponding actions. Formally, we aim to establish a universal policy $\pi$ that predicts future robot observations $o^{r}_{t:t+n}$ and action trajectories $a_{t:t+n}$, leveraging inputs from a sequence of hand demonstration videos $o^{h}_{t:t+n}$, environmental observations $o^{r}_{t}$, and the current proprioception $s_{t}$:
$a_{t:t+n},\ o^{r}_{t:t+n}=\pi(o^{h}_{t:t+n},o^{r}_{t},s_{t}),$ (1)
where $n$ represents the length of the action trajectory we aim to predict. This setting requires the model to not only predict the executed actions but also visually simulate future interactions between the robot arm and objects.
4.1 Architecture
@@
Inspired by the recent success of video diffusion models that are able to predict high-fidelity future frames [20], Human2Robot formulates the behavior learning process from human demonstrations as a generative task—generating robot videos and their corresponding actions simultaneously, conditioned on human videos and historical context. Inspired by  [27], Human2Robot explores a Spatial UNet and a Spatial-Temporal UNet, working collaboratively to learn from humans. In particular, the Spatial UNet extracts features from the robot arm, which is further fed into the Spatial-Temporal UNet for temporal modeling. Human2Robot also contains two behavior extractors that estimate position and motion clues using human videos. To achieve action prediction, we use the intermediate features from the Spatial-Temporal UNet to predict the action trajectory via a Policy Head. Figure 3 gives an overview of our pipeline. Below, we introduce the architecture in detail.
@@
Spatial UNet. The Spatial UNet, consisting of 4 upsampling blocks and 4 downsampling blocks, extracts useful clues from the robot arm. In particular, each block, known as the Spatial Layer, leverages self-attention to learn features tailored for robot arms, as well as cross-attention with pre-extracted CLIP embedding to incorporate semantic clues. The weights of the Spatial-UNet are initialized from those of the Stable Diffusion [48] to ease the learning process. The features derived by Spatial-UNet provide a condition for predicting future frames as well as actions.
@@
While Spatial UNet introduces a similar number of parameters as the denoising UNet, it only extracts features once during the entire process, unlike diffusion-based video generation where each frame undergoes denoising multiple times. Therefore, it does not significantly increase computational overhead during inference.
Behavior Extractor. Recall that we aim to learn from human demonstrations, and thus it is important to extract motion and position information from human videos. This is achieved by Behavior Extractor, which contains four convolution layers (4$\times$ kernels, with a stride of 2). To focus more on the motion trajectories, we employ two independent Behavior Extractors. Behavior Extractor 1 takes the image of the hand corresponding to the robot at the first time step, and Behavior Extractor 2 takes the future image or the entire video of the hand as inputs depending on the training stage, as will be introduced in Section 4.2.
@@
Spatial-Temporal UNet. The Spatial-Temporal UNet aims to predict future frames as well as actions to take simultaneously, requiring the model to learn temporal dynamics. Thus, each block of the Spatial-Temporal UNet (ST-UNet in short) has a Spatial Layer which is the same as the one in Spatial UNet (S-UNet in short), followed by an additional Temporal Layer. More specifically, the ST-UNet takes the features from the behavior extractors and noise latent as inputs. Each Spatial Layer in ST-UNet also concatenates the features of the corresponding layer of the S-UNet, exploring the robot arm as references for prediction. The Temporal Layer focuses on temporal modeling so as to produce high-fidelity future frames as well as accurate action trajectories.
Policy Head. We aim to generate future frames as well as actions to take within a single unified framework. Drawing from the insights of [62], we propose to divide the UNet mapping function into two distinct components: $F=F_{tail}\cdot F_{head}$. We treat $F_{head}$ as a feature extractor for motion information. This is achieved by passing the intermediate features from UNet and proprioception data $s_{t}$ to a light-weighted Multilayer Perceptron to predict the actions of the robotic arm.
@@
Pretrained Encoder and Decoder. Human2Robot builds upon Stable Diffusion [48], which consists of an encoder and a decoder. The encoder turns images into latent embeddings for fast denoising, and the decoder maps the latents back to images. During training, both the encoder and decoder are kept frozen.
4.2 Training
@@
Our approach consists of two training stages. In the first stage, we focus on image generation by taking the first frame of $o^{r}_{t}$, the first frame of human $o^{h}_{t}$ and future frame of human $o^{h}_{t+i}$ as inputs to predict the future frame of robot $o^{r}_{t+i}$. In this stage, we train the S-UNet, the Behavior Extractors, and the ST-UNet without the temporal layers. The spatial layers in both the ST-UNet and S-UNet are instantiated using the powerful open-source Stable Diffusion model (SD)  [48], while the Behavior Extractors are initialized with Gaussian weights, except for the final projection layer, which utilizes zero initialized convolution.
In the second stage, we concentrate the training effort on the temporal layer. In this stage, we take a 30-frame segment of human video $o^{h}_{t:t+n}$ and the first frame $o^{r}_{t}$ of the robot video $o^{r}_{t:t+n}$ as inputs to predict the future robot observation $o^{r}_{t:t+n}$ and action $a_{t:t+n}$. The first frame of the robot video is fed into Behavior Extractor 1, while the entire video segment is passed into Behavior Extractor 2, yielding features $x_{1}^{t}$ and $x_{2}^{t:t+n}$, respectively. Then, we replicate $x_{1}^{t}$ by n times and add it to $x_{2}^{t:t+n}$, which is subsequently added to noise latent and fed to our ST-UNet. The weights of the VAE Encoder and Decoder, as well as the CLIP image encoder are frozen all the time. During our training process, we find that dense sampling (30fps) results in jittering in generated videos. Therefore, we reduce the frame rate to one-fourth of the original, namely cutting the number of frames by three-quarters.
Optimization Objectives. We train Human2Robot with both generation and action prediction objectives, aiming for the model to achieve high-quality video generation and accurate action prediction simultaneously. The goal of generation for the model is to predict the applied noise, so the objective of the optimization process is:
$L_{G}=\mathop{\mathbb{E}_{z_{ts},c,\epsilon,ts}}(||\epsilon-\epsilon_{\theta}(z_{ts},c,ts)||_{2}^{2}),$ (2)
@@
the parameters of the formula represent the latent state $z_{ts}$ at time step $ts$, which is obtained from Spatial-Temporal UNet. The conditioning variable $c$ includes the first frame $o^{r}_{t}$ of the robot and a human video $o^{h}_{t:t+n}$. $\epsilon$ denotes the added noise, and $\epsilon_{\theta}(z_{ts},c,ts)$ represents the model’s predicted noise.
Our action is divided into three components: $a_{t}^{pos}\in\mathbb{R}^{3}$ for 3D position, $a_{t}^{rot}\in\mathbb{R}^{6}$ for rotation and $a_{t}^{grip}\in\{0,1\}$ for gripper state. We use 6D rotation representation proposed in [73, 31] to guarantee continue representation. We use MSE loss $L_{p}$ and $L_{r}$ for position and rotation respectively, and use BCE loss $L_{g}$ for gripper state.
To balance the generation and action prediction losses, we used a set of balancing weights to control the contribution of each loss to the overall objective.
$L=L_{G}+\lambda_{p}L_{p}+\lambda_{r}L_{r}+\lambda_{g}L_{g},$ (3)
where, $\lambda_{p}$, $\lambda_{r}$ and $\lambda_{g}$ are the weights of $L_{p}$, $L_{r}$ and $L_{g}$ respectively.
4.3 Inference
With the aforementioned designs, Human2Robot is able to generate high-quality robot videos and perform actions, using observation captured by the camera of the robot and the provided hand video. We further introduce a set of strategies that facilitate more effective task execution by the robot arm, as will be introduced below.

### v3

@@ Abstract
Distilling knowledge from human demonstrations is a promising way for robots to learn and act. Existing methods, which often rely on coarsely-aligned video pairs, are typically constrained to learning global or task-level features. As a result, they tend to neglect the fine-grained frame-level dynamics required for complex manipulation and generalization to novel tasks. We posit that this limitation stems from a vicious circle of inadequate datasets and the methods they inspire. To break this cycle, we propose a paradigm shift that treats fine-grained human-robot alignment as a conditional video generation problem. To this end, we first introduce H&R, a novel third-person dataset containing 2,600 episodes of precisely synchronized human and robot motions, collected using a VR teleoperation system. We then present Human2Robot, a framework designed to leverage this data. Human2Robot employs a Video Prediction Model to learn a rich and implicit representation of robot dynamics by generating robot videos from human input, which in turn guides a decoupled action decoder. Our real-world experiments demonstrate that this approach not only achieves high performance on seen tasks but also exhibits significant one-shot generalization to novel positions, objects, instances, and even new task categories.
@@ 1 Introduction
Figure 1: Human2Robot: An human-video-conditioned policy, capable of completing seen tasks and one-shot performing unseen tasks with a single human video.
The ability to observe others, whether humans or animals, and acquire skills to solve new tasks is a fundamental reason why humans excel at tackling a wide range of complex challenges. To enable robots to assist with diverse real-world problems, it is essential that they develop a similar capability, particularly the ability to learn directly from human demonstrations. This has motivated extensive research (Bahl et al. 2023; Srirama et al. 2024; Nair et al. 2022; Wang et al. 2023; Bahl, Gupta, and Pathak 2022; Smith et al. 2019) on how to learn from human-provided demonstrations effectively.
Despite recent progress, current approaches face a fundamental generalization gap. While they perform well on tasks encountered during training, they often fail entirely when presented with human demonstrations of unseen tasks. The prevailing paradigm relies on applying self-supervised (Xu et al. 2023b) or contrastive learning (Jain et al. 2024) methods to coarsely aligned human-robot video pairs. However, the lack of fine-grained supervision in existing datasets fundamentally constrains what models can learn. Rather than uncovering detailed action-level correspondences, models are often limited to capturing global features or high-level task summaries. Consequently, many approaches extract holistic video representations using models such as the Perceiver Resampler (Alayrac et al. 2022), which compresses entire clips into fixed-length embeddings. This severely limits their ability to model the nuanced frame-level temporal dynamics that are essential for generalization.
Furthermore, even when densely aligned data are available, models that rely on global feature matching remain fundamentally limited in their ability to capture fine-grained spatio-temporal structure. They often discard the frame-by-frame dynamics that are essential for imitating complex tasks. This limitation has created a vicious circle: the lack of fine-grained datasets gives rise to methods that are incapable of leveraging detailed supervision, while the dominance of such methods in turn discourages the development of fine-grained datasets needed to overcome this barrier. As such, we argue that to achieve true generalization where a robot can perform tasks demonstrated by a human but unseen in its own training data, we must break this cycle. This necessitates a paradigm shift in both learning methodology and data curation.
In this paper, we propose that the key to unlocking fine-grained human-robot alignment lies in video generation (Ho et al. 2022; Esser et al. 2023; Blattmann et al. 2023a; Wu et al. 2023; Khachatryan et al. 2023; Xing et al. 2024; Tu et al. 2024a). Instead of merely learning whether two videos are broadly similar or mapping task goals from the human domain to the robot domain, we aim to generate a corresponding robot video directly from a human demonstration. We posit that by training a model to predict the precise frame-by-frame evolution of a robot’s movements, it learns a far richer and more temporally coherent alignment. This generative process forces the model to internalize the intricate dynamics of manipulation, enabling it to understand how a task is done, not just what is done.
To train such a model would require a dataset with densely aligned and fine-grained human-robot video pairs. Manually annotating such data is prohibitively labor-intensive. To overcome this, we leverage virtual reality teleoperation to create a novel dataset, which we call H&R. By enhancing existing teleoperation systems with improved coordinate system matching, we achieve a seamless mapping between the operator’s hand movements and the robot arm’s motion. This allows us to efficiently collect a large-scale third-person dataset of 2,600 episodes, featuring perfectly synchronized videos of human hands and robot arms across 4 types of basic tasks and 6 long-horizon tasks.
Building on the proposed dataset, we introduce Human2Robot, a framework designed to leverage the proposed H&R dataset. It employs a two-stage training process. First, we train a Video Prediction Model (VPM) built upon a pretrained Stable Diffusion model, which learns to translate a human video into a robot video. This model includes a spatial UNet for feature extraction, a behavior extractors for motion and position encoding, a spatial-temporal UNet that uses a spatial-temporal UNet architecture to explicitly model motion and temporal dynamics, generating a rich latent representation that captures core dynamics of the task. Second, we train an action decoder that conditions on the predictive representations generated by the VPM to output robot actions. This two-stage design effectively learn human-robot alignment and leverages the implicitly learned robot-dynamics features to guide the final policy learning.
With this carefully designed framework, Human2Robot not only excels on seen tasks but also demonstrates remarkable one-shot generalization to novel object positions, appearances, and even entirely new task types and backgrounds. Furthermore, we introduce a KNN-based inference method that allows Human2Robot to perform previously seen tasks with high precision, even without a human demonstration at test time.
@@
  • We present H&R, the first dataset featuring perfectly aligned videos of human hands and robotic arms across a variety of tasks, enabling high-fidelity learning from human demonstrations.
  • We introduce Human2Robot, an end-to-end generative framework utilizes a two-stage training process, which excels in carefully selected tasks, even with variations in positions, appearances, instances, backgrounds and different task types.
  • We propose a KNN+Human2Robot method, which integrates KNN for task prediction, enabling to perform tasks even without human videos as input. This further enhances the scalability and flexibility of the system.
@@ 2 Related Work
Teleoperation. Recently, VR-based methodologies (Iyer et al. 2024; Ding et al. 2024) have attracted considerable attention due to their cost-effectiveness, efficiency and versatility. However, these methods focus on controlling the robot, whereas our goal of using VR is to capture perfectly aligned videos of humans and robots, which are essential for robotic imitation.
Learning from Human Videos. Researchers have recently been attempting to leverage existing human-centric video datasets to enhance robot policy learning (Liu et al. 2018; Smith et al. 2020; Chen, Nair, and Finn 2021; Shaw, Bahl, and Pathak 2022; Zeng et al. 2024). Researchers propose to learn representations from human videos to assist in task execution (Xiao et al. 2022; Wang et al. 2023; Majumdar et al. 2023). However, these approaches need substantial prior knowledge and struggle to transfer to robots (Nair et al. 2022; Bahl et al. 2023; Bahl, Gupta, and Pathak 2022). Meanwhile, some human-video-conditioned methods  (Bharadhwaj et al. 2024; Jain et al. 2024; Xu et al. 2023b) have focused on aligning representations across human and robot videos, they are neither efficient nor capable of generalization. In contrast, our approach uses paired data and diffusion models to achieve strong generalization capabilities.
Diffusion Models for Video Generation. Current research has achieved remarkable performance of video generation (Ho et al. 2022; Esser et al. 2023; Blattmann et al. 2023a; Wu et al. 2023; Khachatryan et al. 2023; Xing et al. 2024; Tu et al. 2024a, b, c). Moreover, diffusion models for video stylization (Ye et al. 2025; Liu et al. 2023) and human image animation (Hu 2024; Tu et al. 2024a) have also shown its promising potential for visual and motion transfer. Moreover, diffusion model not only performs impressively in generative domains but also shows promising applications in visual understanding task. Some notable approaches involve using diffusion models for object detection (Chen et al. 2023), image segmentation (Xu et al. 2023a), and visual representation learning (Yu et al. 2024; Weng et al. 2024). Thus, we aim to utilize the strong capabilities of diffusion model to learn human-robot alignment and leverages the implicitly learned robot-dynamics features to guide the final policy learning.
@@ 3 H&R Dataset
3.1 Coordinate Alignment for Teleoperation
Although teleoperation has become a mainstream method for data collection, no one has yet attempted to use it for collecting paired data of humans and robots. We aim to provide a feasible solution for this.
We identify two core issues. The first issue is that the coordinate system of the robotic arm is not aligned with that of the human hand, which results in a mismatch between the movements of the human hand and the robotic arm in the video. For example, the hand may move across the entire screen, while the robotic arm only moves halfway. This visual difference could lead to difficulties in policy learning. The second issue is the embodiment gap between the human hand and the robotic gripper, which makes teleoperation unsuitable for certain tasks, such as screwing.
To address the issue of coordinate system alignment, we record corresponding three points in the real world of the robotic arm and human hand. We leverage these points to establish a shared coordinate system with consistent scale for both the robotic arm and the human hand. This enables the movement range of the robotic arm to match that of the human hand, as illustrated in Figure 1. Details on how to establish a shared coordinate system can be found in the Supplementary Materials.
However, the embodiment gap remains a challenging issue for current teleoperation system, and we leave it as future work. Therefore, at the data level, we focus on collecting pick-and-place data with human and robotic arm alignment. In our experiments, we evaluate how well the model generalizes after being trained on relatively simple data.
@@ 3.2 Statistics of H&R Dataset
Based on our teleoperation method, we propose our H&R dataset, the first dataset featuring paired human and robot videos. Human events vary from 4 types of basic tasks and 6 long-horizon tasks, as shown in Figure 2. The whole dataset includes 2,600 episodes, and each episode contains frames ranging from 300 to 600. To the best of our knowledge, the H&R dataset is the first video dataset that ensures perfectly aligned video between human and robot.
Figure 2: Dataset Overview. (L) The ratio of four basic task types and long tasks. (R) Platform environment and the object instances used.
@@ 4 Human2Robot
Figure 3: Architecture overview of Human2Robot. Our approach consists of two training stages. In the first stage, we train a Video Prediction Model (VPM) to generate robotic arm videos conditioned on human videos. In the second stage, we freeze the VPM and train an action decoder to predict robot actions based on the motion features generated by the VPM.
In this section, we describe the two-stage training process of our Human2Robot. First, we explore training a VPM to generate robot videos based on human videos, implicitly learning the corresponding robotic arm actions from human movements. Next, we design an action decoder that aggregates the predictive action representations within the VPM and outputs the corresponding robot actions.
4.1 Stage1 : Video Prediction Model (VPM)
@@
We first explored the possibility of generating robot videos from human videos. Inspired by  (Hu 2024), our VPM explores a Spatial UNet (S-UNet in short) and a Spatial-Temporal UNet (ST-UNet in short), working collaboratively to learn from humans. In particular, the S-UNet extracts features from the robot arm, which is further fed into the ST-UNet for temporal modeling. VPM also contains a behavior extractors that estimate position and motion clues from human videos. Figure 3 gives an overview of our pipeline. Below, we introduce the architecture of VPM in detail.
[…]
@@
Spatial UNet. The S-UNet, consisting of 4 upsampling blocks and 4 downsampling blocks, extracts useful clues from the robotic arm. In particular, each block, known as the Spatial Layer (S-Layer in short), leverages self-attention to learn features tailored for robot arms, as well as cross-attention with pre-extracted CLIP embedding to incorporate semantic clues. The weights of the S-UNet are initialized from those of the Stable Diffusion (Rombach et al. 2022) to ease the learning process. The features derived by S-UNet provide a condition for predicting future frames.
@@
While S-UNet introduces a similar number of parameters as the denoising UNet, it only extracts features once during the entire process, unlike diffusion-based video generation where each frame undergoes denoising multiple times. Therefore, it does not significantly increase computational overhead during inference.
@@
Spatial-Temporal UNet. The ST-UNet aims to predict future frames, requiring the model to learn temporal dynamics. Thus, each block of the ST-UNet has a S-Layer which is the same as the one in S-UNet, followed by an additional Temporal Layer (T-Layer in short). More specifically, the ST-UNet takes the features from the behavior extractors and noise latent as inputs. In addition, each S-Layer in ST-UNet also concatenates the features of the corresponding layer of the S-UNet, exploring the robot arm as references for prediction. The T-Layer focuses on temporal modeling so as to produce high-fidelity future frames.
@@
VAE Encoder and Decoder. Human2Robot builds upon Stable Diffusion (Rombach et al. 2022), which consists of an encoder and a decoder. The encoder turns images into latent embeddings for fast denoising, and the decoder maps the latents back to images. During training, both the encoder and decoder are kept frozen.
Training strategy. To generate higher-quality videos, the VPM training process consists of two stages. In the first stage, we focus on image generation to acquire basic generative capabilities. In the second stage, we train it to generate videos.
@@
In the first stage, by taking the first frame of robot $o^{r}_{0}$, the first frame of human $o^{h}_{0}$ and future frame of human $o^{h}_{i}\ (i\in[0,T])$ as inputs to predict the future frame of robot $o^{r}_{i}$, where $T$ refer to the maximum length of video generation. In this stage, we train the S-UNet, the Behavior Extractors, and the ST-UNet without the temporal layers. The S-Layer in both the ST-UNet and S-UNet are instantiated using the powerful open-source Stable Diffusion model (SD)  (Rombach et al. 2022), while the Behavior Extractors are initialized with Gaussian weights, except for the final projection layer, which utilizes zero initialized convolution. In the second stage, we concentrate the training effort on the temporal layer to generate video. In this stage, we take a 30-frame segment of human video $o^{h}_{0:T}$ and the first frame of robot $o^{r}_{0}$ as inputs to predict the future robot observation $o^{r}_{0:T}$. The entire human video segment is first passed into Behavior Extractor, which is subsequently added to noise latent and fed to our ST-UNet. The weights of the VAE Encoder and Decoder, as well as the CLIP image encoder are frozen all the time.
The optimization objectives for this stage are as follows:
[…]
@@
the parameters of the formula represent the latent state $z_{ts}$ at time step $ts$, which is obtained from ST-UNet. The conditioning variable $c$ includes the first frame $o^{r}_{0}$ of the robot and a human video $o^{h}_{0:T}$. $\epsilon$ denotes the added noise, and $\epsilon_{\theta}(z_{ts},c,ts)$ represents the model’s predicted noise.
4.2 Stage2 : Action Decoding.
After the first stage of training, our VPM model is able to visually predict the corresponding robotic arm actions based on human movements. However, fully denoising an entire video is time-consuming, with most of the time spent on reconstructing image-level details, which are unrelated to manipulation. Increasingly, recent research (Hu et al. 2024; Wen et al. 2024; Zhu et al. 2025) has shown that the features of a generative model after a single denoising step already contain sufficient motion features to guide the action head’s planning. Inspired by these works, we treat the pretrained VPM as a vision encoder to extract the predicted robotic arm action information. This means adding intensity-invariant noise (t = K) to the human video and using the first denoised result as the feature. As shown in the Figure 5, we can see that the one-step denoising already contains a lot of action and position information.
Additionally, previous work (Weng et al. 2024; Xiang et al. 2023) has emphasized that the upsampling layers of diffusion models often contain more effective information. Therefore, we use the outputs of the VPM’s upsampling layers as features; more precisely, we use those from the first upsampling layer.
In summary, during this stage, we freeze the parameters of the VPM and treat it as a visual encoder to train the subsequent action decoder. We use the first upsampling layer output $F_{\text{VPM}}$, after the initial denoising step, as the action prior features for the action decoder.
Action Decoder. Inspired by recent video conditioned works (Hu et al. 2024; Luo and Lu 2025), our action decoder consists of two parts: the Video Former (Blattmann et al. 2023b) and the Diffusion Policy (Chi et al. 2023).
The Video Former uses a learnable token $Q$ to aggregate the video features $F_{\text{VPM}}$ into a feature $F_{\text{VF}}\in\mathbb{R}^{n\times L}$ with a fixed-length n and specified channel dimension L. Formally, this branch can be expressed as follows:
$F_{\text{VF}}=\text{FFN}(\text{Temp-Attn}(\text{Spat-Attn}(Q,F_{\text{VPM}}))).$ (2)
We incorporate the aggregated representation $F_{\text{VF}}$ into the diffusion transformer blocks through cross-attention layers. The diffusion policy seeks to reconstruct the original actions $a_{0}$ from the noisy actions $a_{k}=\sqrt{\overline{\beta}_{k}}a_{0}+\sqrt{1-\overline{\beta}_{k}}\epsilon$, where $\epsilon$ represents white noise, and $\overline{\beta_{k}}$ is the noisy coefficient at step $k$. The optimization objectives for diffusion policy are as follows:
[…]
4.3 KNN + Human2Robot.